Debbie is two years older than chuck. The sum of their ages is 30. Find each of their ages.

Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52787)

.
Debbie is two years older than chuck. The sum of their ages is 30. Find each of their ages.
~~~~~~~~~~~~~~~~~~~~

Solution 1. Two equations

D - C =  2,
D + C = 30.

Add the two equations.

Solution 2. One equation.

Let D be the Debbie age. Then the Chuk age is D-2.

We are given that

D + (D-2) = 30.

Simplify and solve.


Answer by Alan3354(69443)

D - C = 2
D + C = 30
--------------- Add